OpenStudy (anonymous):

Which equation shows the substitution method being used to solve the system of linear equations? x – y = 7 x = y + 3 A. (y + 3) – y = 7 B. x – y = y + 3 C. x = (7 – x) + 3 D. x – (y + 3) = 7

OpenStudy (igreen):

We can substitute y + 3 into the first equation for x. \(x - 7 = 3\) Substitute y + 3: \((y + 3) - 7 = 3\) So what's your answer gonna be? @iQueen

OpenStudy (texaschic101):

x - y = 7 x = y + 3......so if we know that x = y + 3, then sub in y + 3 in for x in the first equation (y + 3) - y = 7 any questions ?